OpenStudy (phebe):

How many real number solutions does the equation have? 0=5x^2+2x-12 A- one solution B- two solution C-infinity many solutions D-no solutions

OpenStudy (anonymous):

c bro

OpenStudy (phebe):

are u sure

OpenStudy (anonymous):

yup positive

OpenStudy (phebe):

ok kcool thanx k

OpenStudy (anonymous):

Two solutions were found : x =(-2-√-236)/10=(-1-i√ 59 )/5= -0.2000-1.5362i x =(-2+√-236)/10=(-1+i√ 59 )/5= -0.2000+1.5362i

OpenStudy (helder_edwin):

given the (general) equation \(ax^2+bx+c=0\) then let \(\Delta=b^2-4ac\). If \(\Delta=0\) then 1 real solution. If \(\Delta>0\) then 2 real solutions.

OpenStudy (anonymous):

B

OpenStudy (phebe):

so he was wrong @mjoxprm360

OpenStudy (anonymous):

nope:?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

nope

OpenStudy (phebe):

he told me wrong thanx yall

OpenStudy (anonymous):

1.2 Factoring 5x2 + 2x + 12 The first term is, 5x2 its coefficient is 5 . The middle term is, +2x its coefficient is 2 . The last term, "the constant", is +12 Step-1 : Multiply the coefficient of the first term by the constant 5 • 12 = 60 Step-2 : Find two factors of 60 whose sum equals the coefficient of the middle term, which is 2 .

OpenStudy (anonymous):

The equation will have as many solutions as you have values of x available in the domain. HOWEVER, for one given value of x, there will be ONLY ONE value of y.

OpenStudy (anonymous):

yuppp ^^

OpenStudy (phebe):

soo ok im confused whooosss rit

OpenStudy (anonymous):

Hey add like terms!

OpenStudy (anonymous):

id put c :?

OpenStudy (anonymous):

1.1 Pull out like factors : -5x2 - 2x - 12 = -1 • (5x2 + 2x + 12)

OpenStudy (anonymous):

@phebe its B

OpenStudy (anonymous):

:?

OpenStudy (phebe):

thanx @mjoxprm360 ik ur soo smart thanx

OpenStudy (anonymous):

Anytime :)

OpenStudy (anonymous):

i thought it was b at 1st then c now it might be be:o

OpenStudy (phebe):

XD

OpenStudy (anonymous):

a=5, b=1, c=-12 we get determinant = 1 - 4(5)(-12) = a positive number, therefore, without finding the solutions, we already know it has two real roots.

OpenStudy (anonymous):

..

OpenStudy (phebe):

stop commenting cuz @mjoxprm360 got it rit k

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

neve sed he got wrong but... a=5, b=1, c=-12 we get determinant = 1 - 4(5)(-12) = a positive number, therefore, without finding the solutions, we already know it has two real roots

OpenStudy (anonymous):

@mjoxprm360 u wana help me

OpenStudy (anonymous):

cause immm stuck on 1 :p

OpenStudy (anonymous):

Step-1 : Multiply the coefficient of the first term by the constant 5 • 12 = 60 Step-2 : Find two factors of 60 whose sum equals the coefficient of the middle term, which is 2 . -60 + -1 = -61 -30 + -2 = -32 -20 + -3 = -23 -15 + -4 = -19 -12 + -5 = -17 -10 + -6 = -16 -6 + -10 = -16 -5 + -12 = -17 -4 + -15 = -19 -3 + -20 = -23 -2 + -30 = -32 -1 + -60 = -61 1 + 60 = 61 2 + 30 = 32 3 + 20 = 23 4 + 15 = 19 5 + 12 = 17 6 + 10 = 16 10 + 6 = 16 12 + 5 = 17 15 + 4 = 19 20 + 3 = 23 30 + 2 = 32 60 + 1 = 61 Observation : No two such factors can be found !!

OpenStudy (anonymous):

no not that question//

OpenStudy (anonymous):

http://gyazo.com/1c8dd999bc7ba5f51367a1716ef6b8b2

OpenStudy (anonymous):

this

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

http://gyazo.com/1c8dd999bc7ba5f51367a1716ef6b8b2 is it b?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

thats what i thought thankyou

OpenStudy (anonymous):

no problem.

OpenStudy (anonymous):

i h8 maths so much lol

OpenStudy (anonymous):

I love it :)

OpenStudy (anonymous):

haha will u help me on a question i have no idea about

OpenStudy (phebe):

WHO?

OpenStudy (anonymous):

like im so cluless i dot know where to start

OpenStudy (anonymous):

Absolutly

OpenStudy (anonymous):

thankyou :)

OpenStudy (anonymous):

http://gyazo.com/f368d6f2b97aa68dac39be155869b6bd

OpenStudy (anonymous):

Is there any way you can put that in text format?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The graphs below show four functions. Graph A shows function f of x equals 5 multiplied by 2 to the power of negative x. Graph B shows function f of x equals 5 multiplied by 2 to the power of x. Graph C shows function f of x equals 10 to the power of x. Graph D shows function f of x equals 10 to the power of negative x. Which graph best shows the function f(x) = 5(2)-x to represent the rate at which a radioactive substance decays? Graph A Graph B Graph C Graph D

OpenStudy (anonymous):

ok lets see.

OpenStudy (anonymous):

Would an example help?

OpenStudy (anonymous):

Example 1 The graph of f(x) = x2 is shown at right. For each new function listed below, explain how the new graph would differ from this original graph. g(x) = !2x2 h(x) = (x + 3) 2 j(x) = x2 ! 6 k(x) = 1 4 x2 l(x) = 3(x ! 2) 2 + 7

OpenStudy (anonymous):

graoghing and me is like ... i diont know :(

OpenStudy (anonymous):

lol sorry

OpenStudy (anonymous):

is it a or c xD

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

2 m8s on twitter , 1 said a , 1 said c

OpenStudy (anonymous):

1 said I'm sure that is graph A but don't know how explain when x=0 5(2)^-x = 5*2^0 = 5*1 = 5

OpenStudy (anonymous):

A

OpenStudy (anonymous):

5(2) is 10 : the only graph that goes that high is C : so I'm assuming it is

OpenStudy (anonymous):

ah thankyou :)

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

i like english ;) , im good at that haha

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i gota do 30 questions for homework xD

OpenStudy (anonymous):

im on 18 qand guessesd like 12:p

OpenStudy (anonymous):

lol wow any you need help on?

OpenStudy (anonymous):

would mean alot xD http://gyazo.com/155b55dad19485bea3c19004e44151fb

OpenStudy (anonymous):

its mainly (fx) and i really dont get the fuction and exponents

OpenStudy (anonymous):

what do you think it is?

OpenStudy (anonymous):

your best estimate.

OpenStudy (anonymous):

noo idea would be comlete guess if i said anything xD

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

text format?

OpenStudy (anonymous):

Clara is taking a medicine for a common cold. Her body metabolizes the medicine over a certain time interval. The table below shows the amount of medicine f(t), in mg, that was present in Clara's body after time t. t (hours) 1 2 3 4 5 f(t) (mg) 236.5 223.73 211.65 200.22 189.41 Heidi was administered 300 mg of the same medicine. The amount of medicine in her body f(t) after time t is shown by the equation below. (t) = 300(0.946)t Which statement best describes the rate at which Clara and Heidi's bodies eliminated the medicine? Clara's body eliminated the antibiotic faster than Heidi's body. Clara's body eliminated the antibiotic at the same rate as Heidi's body. Clara's body eliminated the antibiotic at half of the rate at which Heidi's body eliminated the antibiotic. Clara's body eliminated the antibiotic at one-fourth of the rate at which Heidi's body eliminated the antibiotic.

OpenStudy (anonymous):

irst [10:54:09 PM] Alex : Remplace on the second function (to heidi) t by 5 [10:54:21 PM] Alex : and you'll have 227 average [10:54:48 PM] Alex : so they are more antibiotic for heidi [10:55:05 PM] Alex : so clara's antibiotic are eliminate faste

OpenStudy (anonymous):

my friend on skype said this

OpenStudy (anonymous):

for each value in the table, calculate the corresponding value from the function. Now divide one by the other. repeat for the other values in the table. Is that quotient always the same? What does that imply about rate at which the medicine was eliminated?

OpenStudy (anonymous):

o_O

OpenStudy (anonymous):

you want short answer?

OpenStudy (anonymous):

B

OpenStudy (anonymous):

answer is B

OpenStudy (anonymous):

yes ily

OpenStudy (anonymous):

i actually do

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Short answer= answer is B

OpenStudy (anonymous):

ty Which of the following is equal to square root of the cube root of 5? 5 to the power of 1 over 3

OpenStudy (anonymous):

i have never seen a double root:O

OpenStudy (anonymous):

do they exist!!!

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

:o

OpenStudy (anonymous):

hi @phebe

OpenStudy (anonymous):

phebe u know how 2 d it xD